There are 15 ‘True or False’ questions in an examination. In how many ways can a student do the examination if he or she can also choose not to answer almost 3 questions?
Answer by jim_thompson5910(35256)

If the student answers all 15 questions, then there is only one way to do this.

-------------------------------------------------------

If the student skips one question, and exactly one question only, then there are 15 ways to do this (since there are 15 questions to skip).

-------------------------------------------------------

If the student skips two questions, then there are 15 C 2 = 15*14/2 = 105 ways to do this.


-------------------------------------------------------

If the student skips three questions, then there are 15 C 3 = 15*14*13/6 = 455 ways to do this.

=======================================================

So if a student can skip at most 3 questions, then there are 1+15+105+455 = 576 different ways to do this.